The chief ranger of the states Department of Natural Resources is considering a new plan for fighting forest fires in the stFuture Value and Present Value Tables Table 1 Future Value of $1.00(1 + gn Period 4% 6% 8% 10% 12% 14% 20% 1.100 1.210 1.331Period 4% 6% 8% 12% 14% 20% Table II Future Value of a Series of $1.00 Cash Flows (Ordinary Annuity) (1 + 1) - 1 1.000 2.040Table III Present Value of $1.00 (1 + ) Period 1 2 3 26% 794 .630 .500 397 5 .507 189 7 8 9 4% 962 9 25 .889 .855 822 .790 7Table IV Present Value of Series of $1.00 Cash Flows 11. 1 (1 + m) Period 4% 6% 8% 10% 12% 14% 16% 18% 20% 22% 24% 25% 26% 28
Answer #1

Current Plan:

Annual Cost = Staff Salary + Other Cost
= ($200,000 x 8) + ($100,000 x 8) = $2,400,000

Net Present Value = $2,400,000 x 6.145 = $14,748,000

New Proposed Plan:

Initial investment = Building cost + Helicopter & Other equipment + Demolition cost – Salvage value of current equipment
= (200,000 x 4) + (500,000 x 4) + (10,000 x 8) – (60,000 x 8) = $2,400,000

Annual cost = Staff Salary + Other Cost
= (300,000 x 4) + (110,000 x 4) = $1,640,000

Net Present Value = ($1,640,000 x 6.145) – (-2,400,000) = $12,477,800

NPV of incremental cost = NPV of new plan – NPV of current plan
= $12,477,800 -  $14,748,000 = -2,270,200

NPV of incremental cost is negative, the new plan should be selected